2. The bill of materials for product A with the associated component parts and planning factors (in parenthesis) are showed below. How many units of the following components are required to make one unit of the product A? 10 points p. 195 B(2) C (3) ľITRSP-300 H o m ework #3 Fall 2018 (A) Component B (B) Component C (C) Component D (D) Component E (E) Component F (F) Component G
Answer #1

Answer:

a)

Component B :

  • 2 units of component B are required, directly from the chart

b)

Component c:

  • 3 units of component C are required, directly from the chart
  • 1 unit of C per unit of b is also required . So 2 units of C required there.
  • So total 5 units of C per unit of A.

c)

Component d:

  • 1 units of component D is required, directly from the chart
  • 1 unit of D per unit of b is also required . So 2 units of D required there.
  • 1 unit of D per unit of c is also required . So 3 units of D required there.
  • So total 6 units of D per unit of A.

d)

Component E:

  • Each unit of C require 2 units of E
  • So total E required = 2*5 = 10 units (As no of c required = 5)

E)

Component F:

  • Each unit of D require 1 units of F
  • So total F required = 1*6= 6 units (As no of D required = 6)

F)

Component G:

  • Each unit of D require 2 units of G
  • So total G required = 2*6= 12 units (As no of D required = 6)